### v1

@@ Abstract
The rapid development of AI for Science is often hindered by the ”discretization”, where learned representations remain restricted to the specific grids or resolutions used during training. We propose the Neural Proper Orthogonal Decomposition (Neural-POD), a plug-and-play neural operator framework that constructs nonlinear, orthogonal basis functions in infinite-dimensional space using neural networks. Unlike the classical Proper Orthogonal Decomposition (POD), which is limited to linear subspace approximations obtained through singular value decomposition (SVD), Neural-POD formulates basis construction as a sequence of residual minimization problems solved through neural network training. Each basis function is obtained by learning to represent the remaining structure in the data, following a process analogous to Gram–Schmidt orthogonalization. This neural formulation introduces several key advantages over classical POD: it enables optimization in arbitrary norms (e.g., $L^{2}$, $L^{1}$), learns mappings between infinite-dimensional function spaces that is resolution-invariant, generalizes effectively to unseen parameter regimes, and inherently captures nonlinear structures in complex spatiotemporal systems. The resulting basis functions are interpretable, reusable, and enabling integration into both reduced order modeling (ROM) and operator learning frameworks such as deep operator learning (DeepONet). We demonstrate the robustness of Neural-POD with different complex spatiotemporal systems, including the Burgers’ and Navier-Stokes equations. We further show that Neural-POD serves as a high-performance, plug-and-play bridge between classical Galerkin projection and operator learning that enables consistent integration with both projection-based reduced order models and DeepONet frameworks.
@@
Neural Operator, Plug-and-Play, Operator Learning Tool, Reduced Order Modeling, Function Space Orthogonal Bases
1 Introduction
@@
Proper Orthogonal Decomposition (POD), also known as principal component analysis (PCA), is a foundational technique in computational modeling for extracting dominant coherent structures from high-dimensional data. Historically, it emerged from early ideas in statistics and stochastic-process theory (PCA and the Karhunen–Loève viewpoint) and was later adopted in fluid mechanics and turbulence as a dominant way to identify energetic coherent modes from spatiotemporal measurements and simulations [pearson1901liii, hotelling1933analysis, lumley1967similarity, sirovich1987turbulence]. POD supports a broad range of methodologies, including reduced order modeling (ROM), where POD modes provide low-rank representations that enable efficient simulation while preserving essential dynamics [benner2015survey, majda2018strategies, lucia2004reduced, yu2019non], and scientific machine learning (SciML), where POD-based decompositions offer structured low-dimensional representations for learning operators and dynamics (for example, POD–DeepONet architectures [lu2022comprehensive]). Beyond ROM and operator learning, POD is widely used for data compression, feature extraction and modal analysis in complex spatiotemporal systems. Classical POD is typically computed via singular value decomposition (SVD), yielding $L^{2}$-optimal linear basis functions from snapshot data [HLB96, PL98, cazemier1997proper, aubry1993preserving]. Yet, when modern applications demand robustness for different discretizations, diverse regimes and strongly nonlinear behaviours, traditional POD faces persistent obstacles.
@@
These obstacles are well recognized. First, POD is often resolution-dependent: modes learned on one discretization may lose optimality or even effectiveness when the grid changes, limiting transferability across resolutions [taira2017modal, brunton2019data]. Second, POD may lose accuracy even for in-distribution interpolation and can fail more severely for out-of-distribution extrapolation beyond the training regime, which limits its reliability in previously unseen scenarios [carlberg2011efficient, peherstorfer2016data]. Third, although POD is optimal in an $L^{2}$ sense within the training dataset, its linear subspace structure can be too restrictive to represent the nonlinear features that drive many physical systems [quarteroni2015reduced, benner2015survey]. As a consequence, sharp gradients, discontinuities and other strongly nonlinear structures may be poorly captured, especially when such features dominate the dynamics [rowley2004model]. Taken together, these challenges complicate deployment in multi-resolution regimes and hinder real-time simulation, where one needs a compact representation that generalizes reliably as dynamical or physical regimes change.
@@
To address these limitations, we propose neural proper orthogonal decomposition (Neural-POD), a drop-in framework that replaces linear POD modes with nonlinear basis functions parameterized by neural networks. The key idea is to preserve the progressive, mode-by-mode error reduction of POD while upgrading the representational power and enabling transfer across resolutions and regimes. Given snapshot data, Neural-POD learns the first mode by minimizing a reconstruction loss between the snapshots and a neural representation with a learnable time-dependent coefficient; subsequent modes are learned sequentially by training new networks on the residual from the previous approximation [shadden2011lagrangian, yosinski2014transferable, shen2020eeg, deng2021deeplight]. This iterative procedure yields an orthogonal set of Neural-POD modes while directly retaining the structures most relevant to the data. Crucially, after training, only the Neural-POD model parameters (and the associated low-dimensional coefficients) need to be stored, not resolution-specific basis vectors nor the full snapshot matrix, which makes the representation compact and easy to deploy. This compactness enables rapid online evaluation and supports real-time or many-query tasks, where one repeatedly updates reduced states or operators under limited computational budgets.
@@
Neural-POD offers additional benefits that directly target the above challenges. First, the training objective is flexible that allows optimization in task-relevant norms (for example $L^{2}$, $L^{1}$ or $H^{1}$), rather than being prescribed to a fixed $L^{2}$ criterion. Second, representing modes as neural functions admits resolution independence and improves in-distribution and out-of-distribution across multi-resolution settings [mendez2019multi, mendez2020multiscale]. Third, the learned representation can be reused across parameter variations with substantially reduced retraining burden, enabling efficient parametric studies. Finally, Neural-POD can serve as a pretrained, drop-in component inside broader operator-learning frameworks [zahr2015progressive] that provides a structured latent representation that is both compact to store and fast to evaluate for real-time development.
@@
These properties position Neural-POD as a bridge can between projection-based ROM and modern deep learning. In classical ROM, Neural-POD can be inserted into Galerkin projection frameworks to provide a more expressive and transferable basis while retaining the interpretability and structure of projection methods [iliescu2013variational, san2014proper, KV01, iliescu2014are, mou2021data]. In operator learning, Neural-POD offers a principle low-dimensional representation that can improve both efficiency and interpretability. For instance, Deep Operator Networks (DeepONets) comprise a branch network that encodes the input function sampled at sensor locations and a trunk network that encodes query locations for the output [lu2019deeponet, wang2021learning, winovich2025active, lu2025evolutionary, goswami2023physics]. By using Neural-POD as the branch network, DeepONets can incorporate orthogonal, data-driven nonlinear modes that capture underlying dynamics while remaining resolution-independent. This integration can improve generalization across discretizations and reduce training costs (since only Neural-POD can be pretrained), thereby enabling more efficient and robust operator learning in real-time or near-real-time settings. In brief summary, our primary contributions include the following:
@@
  • Neural-POD is a plug-and-play component for both Galerkin ROM and operator learning: once the basis/model is learned offline, it can be dropped into a reduced order model with minimal modification, and the online stage only requires assembling and evolving the reduced system for fast, easy deployment across parameterized PDEs (see Figure 1(b)). This bridges traditional ROM and surrogate modeling with operator learning by fitting naturally into familiar offline/online workflows and existing Galerkin solvers (see Figure 3). Moreover, the same plug-and-play interface enables straightforward reuse across PDEs with different parameters which supports a pretrained, open-source database and a shared repository of benchmark problems spanning diverse geometries and parameter settings.
@@
  • Neural-POD also has strong educational value as an easy-to-use ROM and operator-learning tool. (see Figure 4) Its modular, plug-and-play property makes it well-suited for classroom adoption (e.g., undergraduate numerical analysis, scientific computing, and introductory SciML courses), where students can rapidly learn reduced order models and operator learning, explore the offline/online procedures, and build intuition for projection-based modeling without heavy software overhead. This positions Neural-POD as a practical foundation for course modules, hands-on assignments, and open teaching resources with broad educational impact.
@@
  • Neural-POD enables accurate in-distribution prediction and reasonable generalization for out of distribution prediction by learning resolution-independent basis functions that remain effective under modest parameter shifts, without retraining (see Figure 10 and Table 1).
@@
  • Neural-POD constructs nonlinear, orthogonal basis functions through an iterative residual minimization procedure using with neural networks that is analogous to Gram–Schmidt orthogonalization and capable of capturing complex features (see Figure 2).
@@
  • Neural-POD supports training under different norms (e.g., $L^{2}$, $L^{1}$), which enables the learned bases to reflect smoothness, discontinuities, or other structural characteristics induced by the chosen norm (see Figures 79).
@@
The above novelties are also illustrated in Figures 1(a) and 1(b). The remainder of this paper is organized as follows. Section 2 provides a detailed description of the Neural-POD architecture and associated algorithms, and shows how Neural-POD can be incorporated into both Galerkin projection-based reduced order models and DeepONet operator learning frameworks. Section 3 shows the numerical results on Burgers’ equation and the two‐dimensional Navier–Stokes flow. Finally, Section 4 summarizes the conclusions and outlines directions for future work.
@@
Figure 1: Neural-POD connects classical projection-based model reduction and modern operator learning. By representing basis functions as neural networks, Neural-POD enables resolution-independent, parameter-aware ROMs and can be integrated with DeepONet operator learning frameworks.
2 Neural Proper Orthogonal Decomposition (Neural-POD)
2.1 Classic Proper Orthogonal Decomposition (POD)

### v2

@@ Abstract
AI for science (AI4Science) models often suffer from discretization: learned representations remain tied to the training grid, limiting transfer across resolutions, solvers and applications. We introduce Neural Proper Orthogonal Decomposition (Neural-POD), a plug-and-play neural operator that learns nonlinear, orthogonal basis functions directly in function space and can be integrated in both projection-based reduced order models and operator-learning frameworks such as DeepONet. Neural-POD replaces SVD-derived, resolution-dependent linear modes with continuous, resolution-invariant bases learned via sequential residual minimization, analogous to Gram–Schmidt orthogonalization. The framework supports training under task-specific norms (e.g., $L^{2}$, $L^{1}$), improves out-of-distribution generalization to unseen parameter regimes, and captures nonlinear structure in complex systems. Because the learned bases are interpretable and reusable, Neural-POD serves as a general representation module for AI4Science workflows. We demonstrate Neural-POD on Burgers’ and Navier–Stokes equations.
@@
Neural Operator, Plug-and-Play, Operator Learning, Reduced Order Modeling, Function Space Orthogonal Bases
@@
AI for science (AI4Science) increasingly relies on learned representations that must transfer reliably across discretizations, geometries, and parameter regimes, rather than being tied to a particular grid or snapshot resolution. In this setting, a central challenge is to build low-dimensional structures that are both faithful to the underlying infinite-dimensional dynamics and reusable across solvers and data that supports efficient simulation without retraining from scratch whenever the discretization changes. Consequently, Proper Orthogonal Decomposition (POD) is widely used in AI4Science as a physics-informed representation tool that provides compact low-dimensional bases for high-dimensional data that support efficient learning, inference and prediction across different scientific tasks. POD, also known as principal component analysis (PCA), is a foundational technique in computational modeling for extracting dominant coherent structures from high-dimensional data. Historically, it emerged from early ideas in statistics and stochastic-process theory (PCA and the Karhunen–Loève viewpoint) and was later adopted in fluid mechanics and turbulence as a dominant way to identify energetic coherent modes from spatiotemporal measurements and simulations [pearson1901liii, hotelling1933analysis, lumley1967similarity]. POD supports a broad range of methodologies, including reduced order modeling (ROM), where POD modes provide low-rank representations that enable efficient simulation while preserving essential dynamics [benner2015survey, lucia2004reduced, yu2019non], and scientific machine learning (SciML), where POD-based decompositions offer structured low-dimensional representations for learning operators and dynamics (for example, POD–DeepONet architectures [lu2022comprehensive]). Beyond ROM and operator learning, POD is widely used for data compression, feature extraction and modal analysis in complex spatiotemporal systems. Classical POD is typically computed via singular value decomposition (SVD), yielding $L^{2}$-optimal linear basis functions from snapshot data [HLB96]. Yet, when modern applications demand robustness for different discretizations, diverse regimes and strongly nonlinear behaviours, traditional POD faces persistent obstacles.
@@
These obstacles are well recognized. First, POD is often resolution-dependent: modes learned on one discretization may lose optimality or even effectiveness when the grid changes, limiting transferability across resolutions [taira2017modal]. Second, POD may lose accuracy even for in-distribution interpolation and can fail more severely for out-of-distribution extrapolation beyond the training regime, which limits its reliability in previously unseen scenarios [carlberg2011efficient]. Third, although POD is optimal in an $L^{2}$ sense within the training dataset, its linear subspace structure can be too restrictive to represent the nonlinear features that drive many physical systems [quarteroni2015reduced, benner2015survey]. As a consequence, sharp gradients, discontinuities and other strongly nonlinear structures may be poorly captured, especially when such features dominate the dynamics [rowley2004model]. Taken together, these challenges complicate deployment in multi-resolution regimes and hinder real-time simulation, where one needs a compact representation that generalizes reliably as dynamical or physical regimes change.
@@
To address these limitations, we propose neural proper orthogonal decomposition (Neural-POD), a drop-in framework that replaces linear POD modes with nonlinear basis functions parameterized by neural networks. The key idea is to preserve the progressive, mode-by-mode error reduction of POD while upgrading the representational power and enabling transfer across resolutions and regimes. Given snapshot data, Neural-POD learns the first mode by minimizing a reconstruction loss between the snapshots and a neural representation with a learnable time-dependent coefficient; subsequent modes are learned sequentially by training new networks on the residual from the previous approximation [shadden2011lagrangian, yosinski2014transferable, deng2021deeplight]. This iterative procedure yields an orthogonal set of Neural-POD modes while directly retaining the structures most relevant to the data. Crucially, after training, only the Neural-POD model parameters (and the associated low-dimensional coefficients) need to be stored, not resolution-specific basis vectors nor the full snapshot matrix, which makes the representation compact and easy to deploy. This compactness enables rapid online evaluation and supports real-time or many-query tasks, where one repeatedly updates reduced states or operators under limited computational budgets.
@@
Neural-POD offers additional benefits that directly target the above challenges. First, the training objective is flexible that allows optimization in task-relevant norms (for example $L^{2}$, $L^{1}$ or $H^{1}$), rather than being prescribed to a fixed $L^{2}$ criterion. Second, representing modes as neural functions admits resolution independence and improves in-distribution and out-of-distribution across multi-resolution settings [mendez2019multi]. Third, the learned representation can be reused across parameter variations with substantially reduced retraining burden, enabling efficient parametric studies. Finally, Neural-POD can serve as a pretrained, drop-in component inside broader operator-learning frameworks [zahr2015progressive] that provides a structured latent representation that is both compact to store and fast to evaluate for real-time development.
@@
These properties position Neural-POD as a bridge can between projection-based ROM and modern deep learning. In classical ROM, Neural-POD can be inserted into Galerkin projection frameworks to provide a more expressive and transferable basis while retaining the interpretability and structure of projection methods [KV01, iliescu2014are, mou2021data]. In operator learning, Neural-POD offers a principle low-dimensional representation that can improve both efficiency and interpretability. For instance, Deep Operator Networks (DeepONets) comprise a branch network that encodes the input function sampled at sensor locations and a trunk network that encodes query locations for the output [lu2019deeponet, winovich2025active, lu2025evolutionary]. By using Neural-POD as the branch network, DeepONets can incorporate orthogonal, data-driven nonlinear modes that capture underlying dynamics while remaining resolution-independent. This integration can improve generalization across discretizations and reduce training costs (since only Neural-POD can be pretrained), thereby enabling more efficient and robust operator learning in real-time or near-real-time settings. In brief summary, our primary contributions include the following:
@@
  • Neural-POD is a plug-and-play component for both Galerkin ROM and operator learning: once the basis/model is learned offline, it can be dropped into a reduced order model with minimal modification, and the online stage only requires assembling and evolving the reduced system for fast, easy deployment across parameterized PDEs (see Figure 1(b)). This bridges traditional ROM and surrogate modeling with operator learning by fitting naturally into familiar offline/online workflows and existing Galerkin solvers (see Figure 2(a)). Moreover, the same plug-and-play interface enables straightforward reuse across PDEs with different parameters which supports a pretrained, open-source database and a shared repository of benchmark problems spanning diverse geometries and parameter settings.
@@
  • Neural-POD also has strong educational value as an easy-to-use ROM and operator-learning tool. (see Figure 2(b)) Its modular, plug-and-play property makes it well-suited for classroom adoption (e.g., undergraduate numerical analysis, scientific computing, and introductory SciML courses), where students can rapidly learn reduced order models and operator learning, explore the offline/online procedures, and build intuition for projection-based modeling without heavy software overhead. This positions Neural-POD as a practical foundation for course modules, hands-on assignments, and open teaching resources with broad educational impact.
@@
  • Neural-POD enables accurate in-distribution prediction and reasonable generalization for out of distribution prediction by learning resolution-independent basis functions that remain effective under modest parameter shifts, without retraining (see Figure 4(d)4(e) and Table 1).
@@
  • Neural-POD constructs nonlinear, orthogonal basis functions through an iterative residual minimization procedure using with neural networks that is analogous to Gram–Schmidt orthogonalization and capable of capturing complex features (see Figure 1(c)).
@@
  • Neural-POD supports training under different norms (e.g., $L^{2}$, $L^{1}$), which enables the learned bases to reflect smoothness, discontinuities, or other structural characteristics induced by the chosen norm (see Figures 4(a)4(c)).
@@
The above novelties are also illustrated in Figures 1(a) and 1(b).
Rapid progress in AI4Science for reduced representations and operator learning has exposed a key gap: performance can be highly sensitive to discretization choices (resolution, numerical schemes, and discretization error), limiting reliability and generalization for different regimes. We address this by studying the robustness of Neural-POD’s nonlinear functional representations and integrating them into (i) projection-based ROMs and (ii) DeepONet-style operator learning. We further assess performance under a fixed training budget to identify when Neural-POD remains effective. Using Burgers’ equation and 2D Navier–Stokes, we benchmark Neural-POD against classical POD in reconstruction, predictive accuracy, and robustness to in-distribution and out-of-distribution parameters.
@@
(c) Overview of Neural-POD training. Snapshots $u(x,t,\kappa)$ are collected over time for varying parameter values $\kappa$ (upper left). The training architecture decomposes each snapshot into products of spatial mode functions $\Phi_{i}(x;\kappa)$ and temporal coefficients $\Psi_{i}(t;\kappa)$, which guarantees the orthogonality for all modes (bottom left). Once trained, the Neural-POD basis provides plug-and-play reduced representations that can be evaluated on any discretization $\hat{x}$ and parameter choice $\hat{\kappa}$ (right, $\hat{\kappa}$ may differ from the training snapshots’ parameter $\kappa$).
Figure 1: Neural-POD for resolution-independent model reduction and operator learning.
(a) Schematic diagram of the proposed Neural-POD-ROMs and traditional POD-based reduced order models (POD-ROMs). Top portion: the Neural-POD extension that enables plug-and-play operation on arbitrary meshes and at arbitrary resolution, informed by pretrained neural representations of POD bases and PDE parameters. Bottom portion: the traditional POD flowchart, including snapshot generation for parameter $\kappa$, computation of full POD modes via SVD, construction of the reduced basis, and Galerkin projection of the governing PDE.
(b) Schematic of the proposed Neural-POD Deep Operator Network (NP-DeepONet). The branch network takes initial condition fields as input, while the trunk network evaluates functions at spatial query points. A pretrained Neural-POD module provides reduced basis information $\phi_{1},\phi_{2},\ldots,\phi_{P}$} obtained from snapshot data. The outputs are combined to produce the operator evaluation $G_{\theta}(u_{m})(x)$, enabling efficient and data-driven surrogate modeling of PDE solutions in reduced coordinates.
Figure 2: Neural-POD integrations for reduced order modeling and operator learning.
Results
Neural-POD for Burgers’ equation
$u(x,0)=u_{0}(x),$ $x\in(0,1),$ (2)
In the following experiments, the Neural-POD model is trained using snapshot data generated at viscosity values $\nu$ sampled within the interval $[0.005,,0.01]$. This parameter range defines the training domain for the Neural-POD basis learning. Figure 3(a) shows the evolution of the one-dimensional Burgers equation for different sampled viscosity values $\nu$ used to construct the training dataset for Neural-POD. As $\nu$ decreases (right to left), the solution transitions from a smooth diffusive profile to one dominated by advective nonlinearities, leading to the formation of sharp fronts and nearly discontinuous shock structures. This coupled effects of convection and diffusion underscores the limits of traditional POD in handling unseen parameters, motivating parametric POD formulations that adapt to viscosity and smoothness variations. Extended Data Figure 1 shows the first three Neural-POD modes trained from FOM data for different viscosity values. As the viscosity changes, the Neural-POD modes behaves significantly different behavior, with $\nu$ decreases, Neural-POD becomes increasingly oscillatory and sharply localized near the shock region. Such sensitivity of the Neural-POD basis to the viscosity parameter indicates that the modes are not easily transferable across regimes, making it challenging to construct a compact parametric traditional POD representation that remains accurate for a wide range of different viscosity flows.
Neural-POD with $L^{1}$ and $L^{2}$ loss
Extended Data Figure 2 illustrates the reconstruction of spatiotemporal fields using the Neural-POD model for a viscosity case not included in the training data. The results are shown for latent dimensions of 2, 4, and 6. Since the traditional POD basis depends on snapshot data from the training parameters, it cannot be constructed for this unseen case. In contrast, the Neural-POD can accurately reconstruct the spatio-temporal dynamics even with only two or four modes which shows strong generalization capability and the ability to capture nonlinear correlations between spatial and temporal features beyond the training regime. As the dimension of modes increases, the Neural-POD reconstructions further improve and closely approach the truth solution, consistently yielding lower reconstruction errors and better preservation of localized fine-scale structures over time. For the case with $r=6$, the model trained with the $L^{1}$ loss exhibits better performance compared to its $L^{2}$ loss Neural-POD, as it more effectively captures the steep gradients and localized features. Figure 3(b) compares the first three spatial modes obtained from the classical POD and those learned through Neural-POD training using $L^{1}$- and $L^{2}$-based losses across different viscosity regimes. The $L^{2}$-optimal Neural-POD tends to emphasize smooth, energy-dominant structures, while the $L^{1}$-based formulation better preserves localized sharp transitions by penalizing large pointwise deviations less severely. Figure 4(a) shows that the Neural-POD residual decreases rapidly with the number of modes, closely mirroring the decay of the classical POD residual.
(a) Solution snapshots of the Burgers equation at different sampled viscosity values $\nu$ used for Neural-POD training.
(b) First three traditional POD modes and Neural-POD under different loss choices, i.e., $L^{1}$ and $L^{2}$ norms, for different viscosity values.
Figure 3: Burgers equations: (a) solution snapshots for different viscosities $\nu$; (b) traditional POD modes versus Neural-POD bases learned with $L^{1}$ and $L^{2}$ losses.
(a) Neural-POD residuals versus modes $r$ for the $L^{2}$ loss (left) and the $L^{1}$ loss (right).
(b) $L^{2}$ norm of the residuals for $L^{1}$- and $L^{2}$-optimal Neural-POD reconstructions.
(c) $L^{1}$ norm of the residuals for $L^{1}$- and $L^{2}$-optimal Neural-POD reconstructions.
(d) In-distribution results of the first three POD modes: comparison between traditional POD with extra snapshots and Neural-POD for an unseen viscosity value $\nu$.
(e) Out-of-distribution results of the first three POD modes: comparison between traditional POD with extra snapshots and Neural-POD for an unseen viscosity value $\nu$.
Figure 4: Summary of Burgers Neural-POD results: (a) training residual decay versus mode index for $L^{2}$ and $L^{1}$ losses; (b–c) reconstruction residuals under $L^{2}$ and $L^{1}$ metrics across representative viscosities; (d–e) in-distribution and out-of-distribution comparisons of POD (with extra snapshots) and Neural-POD for an unseen viscosity $\nu$.
Figures 4(b)4(c) compares the $L^{1}$ and $L^{2}$ norms of the reconstruction residuals obtained from Neural-POD models trained with $L^{1}$- and $L^{2}$-based loss functions across different viscosity values. Panels 4(b) and 4(c) correspond to the $L^{2}$ and $L^{1}$ residual norms, respectively, evaluated for three representative cases of $\nu=0.01$, $\nu=0.03$, and $\nu=0.05$. In panel 4(b), the $L^{2}$-optimal Neural-POD produces smooth and globally accurate reconstructions, with residuals remaining low throughout the spatial domain, particularly for higher viscosity values where the flow is predominantly diffusive. Panel 4(c) presents the corresponding $L^{1}$ residual norms for both the $L^{1}$- and $L^{2}$-trained models. The $L^{1}$-optimal Neural-POD exhibits improved accuracy in certain cases, particularly for specific reduced dimensions $r$ ($r=1$ and $r=3$) where it better captures localized steep gradients and nonlinear features. Because the $L^{1}$ norm penalizes large deviations less severely, it allows for sharper reconstructions in regions of strong nonlinearity, while the $L^{2}$ formulation maintains smoother but more globally averaged behavior. Overall, the comparison indicates that the relative performance of the two loss formulations depends on both the viscosity parameter and the dimension $r$.
Neural-POD for in-distribution and out-of-distribution viscosity values
In the following experiments, the Neural-POD model is trained under the same numerical setting as in Figure 3(a), using snapshot data generated at viscosity values $\nu$ sampled within the interval $[0.005,,0.01]$. This interval defines the training parameter domain for Neural-POD basis learning. Figures 4(d)4(e) show both the in-distribution and out-of-distribution performance of the Neural-POD framework. Here, in-distribution refers to evaluating the learned Neural-POD model at viscosity values $\nu$ within the parameter interval used for training, while out-of-distribution refers to evaluating the model at $\nu$ values outside that interval. More specifically, Figure 4(d) and Figure 4(e) compare the first three basis modes from traditional POD with those produced by Neural-POD trained using the $L^{2}$ loss. Because traditional POD constructs modes only at the discrete viscosity levels contained in the snapshot data, it cannot directly produce basis functions for intermediate or out-of-range parameter values. Neural-POD, on the other hand, learns a continuous mapping from $\nu$ to the modal space, allowing it to generate basis functions for both unseen in-distribution and out-of-distribution parameter regimes. Notably, even for the fully unseen parameter value $\nu=4\times 10^{-3}$ (see Figure 4(e)), which lies outside the training interval $[0.005,,0.01]$ and represents a $20\%$ out-of-distribution shift below its lower bound, Neural-POD still generates modes that remain consistent with the corresponding POD modes that confirms its generalization capability.
Neural-POD for Navier-Stokes equations
$\nabla\cdot u=0\,$, (7)
where $u$ is the velocity, $p$ the pressure, and $Re$ the Reynolds number. The computational domain is a $2.2\times 0.41$ rectangular channel with a radius =0.05=0.05 cylinder, centered at $(0.2,0.2)$, see Figures 6(a)6(b). We prescribe no-slip boundary conditions on the walls and cylinder, and the following inflow and outflow profiles:
$u_{1}(0,y,t)$ $=u_{1}(2.2,y,t)=\frac{6}{0.41^{2}}y(0.41-y),$ (8)
$u_{2}(0,y,t)$ $=u_{2}(2.2,y,t)=0,$ (9)
where $u=\langle u_{1},u_{2}\rangle$.
The dataset used is generated by a high-fidelity finite element simulation of the unsteady Navier–Stokes equations at $Re=50$. The spatial discretization employs a Taylor–Hood element pair (quadratic velocity, linear pressure) on an unstructured triangular mesh that provides sufficient resolution of the boundary layers and wake structures. The resulting finite element discretization yields approximately $1.2\times 10^{4}$ velocity degrees of freedom (DOFs). Temporal integration is performed using a second-order backward differentiation formula (BDF2). A total of $1000$ velocity snapshots are recorded uniformly in time over the interval $[10,20]$. To make a consistent comparison, these snapshots are used to generate both the Proper Orthogonal Decomposition (POD) basis and to train the Neural–POD. Figures 6(c)6(d) illustrates the first two spatial modes identified by the classical POD and Neural–POD. The close agreement between the two confirms the consistency of Neural–POD with the traditional POD framework.
Neural-POD in Reduced Order Modeling
[…]
and $\nu=10^{-3}$. This test problem has been used in [mou2021data, koc2023uniform].
Snapshot Generation
The truth data are obtained using a linear finite element spatial discretization with mesh size $h=1/512$ and a Crank–Nicolson time discretization with timestep size $\Delta t=10^{-4}$. A total of $101$ snapshots are collected at equally spaced time instances over the interval $t\in[0,1]$ for constructing the POD and Neural-POD modes. More details are provided in Supplementary Section A.2. and Supplementary Data, Fig. A3.
(a) Geometry of two-dimensional flow past the cylinder.
(b) Sampling points for the flow past the cylinder test problem.
(c) First two POD modes.
(d) First two Neural-POD modes.
Figure 6: (a–b) Geometry and sampling points of the two-dimensional flow past a cylinder problem. (c–d) First two modes of POD and Neural-POD for two-dimensional flow past the cylinder with $Re=50$.
Discussion of Numerical Results
Figure 3 compares the modes from classical POD and Neural–POD under different optimality criteria for Burgers equation with initial condition in (15) . The $L^{2}$-optimal Neural–POD modes closely resemble the classical POD modes, showing that the neural representation preserves the same energy-optimization structure when trained in the same norm. In contrast, the $L^{1}$-optimal Neural–POD produces sharper and more localized modes, emphasizing regions with strong gradients or discontinuities. This indicates that while the $L^{2}$-trained Neural–POD recovers the traditional POD basis, the $L^{1}$ formulation provides a flexible alternative for capturing nonlinear and nonsmooth dynamics. We consider two projection-based ROMs: the standard POD-ROM, which uses classical POD modes, and the Neural-POD-ROM, which employs modes learned via Neural-POD.
$\nu$ Model $r=2$ $r=4$ $r=6$
0.002 POD–ROM 2.10e-01 1.66e-01 1.76e-01
Neural–POD–ROM ($L^{2}$) 2.11e-01 1.67e-01 1.77e-01
0.0025 POD–ROM No in distribution ability
Neural–POD–ROM ($L^{2}$) 9.61e-02 1.18e-01 1.35e-01
0.005 POD–ROM No out of distribution ability
Neural–POD–ROM ($L^{2}$) 1.91e-01 1.21e-01 8.46e-02
[…]
Neural-POD in Deep Operator Learning
We consider the one-dimensional Burgers’ equations (1)–(3) following the same configuration described in Section Results. The dataset is adopted from [lu2022comprehensive], where the initial conditions are sampled from a Gaussian random field characterized by a Riesz kernel, $\mu=\mathcal{R}(0,,625(-\Delta+25I)^{-2})$, with $\mu$ denoting the probability measure over the function space and $\Delta$ and $I$ representing the Laplacian and identity operators, respectively. Both input and output fields are discretized on a uniform grid with 128 spatial points.
In contrast to the method proposed in [lu2022comprehensive], which employs snapshots at $t=1$ to construct the POD basis, our Neural-POD framework learns the basis functions directly from the temporal evolution of the Burgers’ equation across multiple time steps. We employ four latent modes in the reduced representation and train the model on 128-point spatial grids, while evaluating its reconstruction and generalization performance on coarser 64-point grids. This setup enables assessment of both the model’s spatial resolution robustness and its capacity to generalize across varying discretizations and viscosity parameters. This is further confirmed by Extended Data Tables 1. At the coarse resolution ($N=64$), the Neural-POD–DeepONet achieves lower testing errors than the traditional POD–DeepONet, with roughly a $20\%$ reduction in both $L_{1}$ and $L_{2}$ metrics. These results, consistent with Extended Data Figure 2(b), show that the Neural-POD framework offers clear benefits under limited resolution, while recovering classical POD behavior as the discretization becomes sufficiently fine. At coarse resolution ($N=32$), the Neural-POD–DeepONet exhibits noticeably improved reconstructions over the traditional POD–DeepONet, capturing sharper transitions. This improvement can be attributed to the adaptive and data-driven nature of the Neural-POD basis, which enhances the representation of localized features under limited spatial resolution. As the training resolution increases to $N=64$, the difference between the two approaches becomes less pronounced, though the Neural-POD still achieves slightly lower reconstruction errors. For the finest resolution ($N=128$), both Neural-POD–DeepONet and POD–DeepONet produce nearly similar results, indicating that the advantage of the learned basis diminishes as the resolution becomes sufficiently high to resolve the dominant flow structures. In this regime, the $L^{1}$-based Neural-POD performs marginally worse than the $L^{2}$ variant due to its less stringent global smoothness constraint. Overall, the results suggest that the Neural-POD framework provides clear advantages at coarse or moderately resolved training settings, where it compensates for the loss of spatial detail through adaptive modal learning, while converging to the classical POD behavior at higher resolutions. Extended Data Figure 4 shows the comparison between the proposed Neural-POD–DeepONet and the traditional POD–DeepONet with different spatial training resolutions, $N=32$, $N=64$, and $N=128$. Each model is evaluated on unseen initial conditions of the Burgers’ equation to verify its ability to generalize for both data realizations and discretization levels.
Discusssions
Future research will focus on the following different directions: First, we will extend the Neural-POD framework to more challenging and high-dimensional scientific problems such the turbulent three-dimensional Navier–Stokes or Boussinesq equations, where nonlinear modal interactions and multi-scale energy cascades provide a demanding test [mou2023energy, bright2013compressive, lin2009efficient]. Another promising direction is the integration of Neural-POD with physics-informed neural networks (PINNs) to enforce physical constraints during basis construction, thereby improving interpretability and extrapolation for physics-guided model reduction [cai2021physics, lu2025mopinnenkf]. Furthermore, coupling Neural-POD with data assimilation techniques such as Ensemble Kalman Filters (EnKF) or variational methods may enable real-time state estimation and model correction in dynamical systems [popov2021multifidelity, bilionis2013multi]. Also, in digital-twin applications, Neural-POD can be employed to build efficient and adaptive surrogate models that enable real-time monitoring, prediction, and control of complex physical systems [xiu2025computational, jones2020characterising]. Lastly, we plan to extend the framework to parameterize and learn geometrical variations that enables the community to contribute new mesh configurations and expand the applicability of Neural-POD to complex domains with heterogeneous geometries [shukla2024deep].
Methods
Classic Proper Orthogonal Decomposition (POD)